Which expression shows that 11 + 7 = 5 + 13?
A.
(5 + 6) + 7 = 5 + (6 + 7)
B.
6 + 6 + 6 = 18
C.
5 + 9 + 4 = 6 + 7 + 5
D.
(11 + 4) + 3 = (5 + 7) + 6

Found 2 solutions by oscargut, stanbon:
Answer by oscargut(2103)

Answer: A
If you need more help with similar questions you can contact me now at:
mthman@gmail.com
Thanks

Answer by stanbon(75887)

Which expression shows that 11 + 7 = 5 + 13?
Answer:::
(5 + 6) + 7 = 5 + (6 + 7)
=========
Cheers,
Stan H.
=========